x^4-9x^2+20 Can you please help me solve this quadratic equation
Found 2 solutions by Alan3354, ReadingBoosters:
Answer by Alan3354(69443)

x^4-9x^2+20 Can you please help me solve this quadratic equation
--------------
There's no equal sign, so it's not an equation.

Answer by ReadingBoosters(3246)

Starting with
x^4 - 9x^2 + 20 = 0
(x^2 - )(x^2 - ) = 0 both negative due to - 9 and + 20
Find common multiple of 20 that adds to -9
(x^2-5)(x^2-4) = 0
(x^2-4) can be factored further
(x^2-5)(x+2)(x-2)
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com